OpenStudy (anonymous):

Which of the following statements is not correct? 0.20 mole of O2= 6.4 g 0.75 mole of H2CO3= 47 g 3.42 moles CO = 95.8 g 4.1 moles Li2O = 94 g I need help please :/

OpenStudy (anonymous):

Well this isn't chemistry here But still: http://www.lenntech.com/calculators/molecular/molecular-weight-calculator.htm In here I found that molar mass of Oxygen (O) is about 16g/mol and for Lithium (Li) it is about 7 so the 4th one goes: \[ Li2O_{g/mol} = Li_{g/mol} + 2 \cdot O_{g/mol} = 7_{g/mol} + 32_{g/mol} = 39_{g/mol} \\ Li2O_{g/mol} \cdot 4.1_{mol} = 39_{g/mol} \cdot 4.1_{mol} = 159.9_g \\ 159.9_g \ne 94_g \] Hope I got it right =S